OpenStudy (anonymous):

Jack was 9 years older than Priscilla. Together their ages totaled 165 years. What was Jack's age?

OpenStudy (conqueror):

165 - 9 = 156 / 2 = 78.

OpenStudy (conqueror):

Do you understand? @scoutuj

OpenStudy (anonymous):

So the total age minus the 9 yeras and then divide it by 2

OpenStudy (anonymous):

78

OpenStudy (anonymous):

its 78

OpenStudy (anonymous):

ok

OpenStudy (conqueror):

Yes, divide it by 2 people.

OpenStudy (anonymous):

@Conqueror

OpenStudy (conqueror):

No wait, that's Priscilla's age, Add 78 + 9 and that's your answer.

OpenStudy (anonymous):

okay it said the answer was 87 not 78... so??

OpenStudy (anonymous):

could you help me on mine

OpenStudy (anonymous):

87

OpenStudy (anonymous):

okay well here is one similar... Jack is 10 years older than Priscilla. Together their ages totaled 170 years. whats Jacks age? So how would i figure it out

OpenStudy (conqueror):

Subtract total age by the total amount of years older Jack is, then divide the result by 2 and that's your answer.

OpenStudy (conqueror):

No wait, add the second result with the total amount of years.

OpenStudy (conqueror):

170 - 10 = 160 / 2 = Priscilla's age + 10 = Jack's age. There, a lot less confusing.

OpenStudy (anonymous):

thank you